A person travels 600 km to his home , partly by train and partly by car . He takes 8 hours of he travels 120 km by train and the rest by car . Find the speed of train and the car ?
Answer by stanbon(75887)

Train Data:
distance=120km; time=x hrs;
Since rate=distance/time, rate=120/x
Car Data:
distance=600-120=480km; time=8-x; rate=?480/(8-x)
EQUATION:
There is not enough information to write an equation.
Check again the statement of this problem in the text.
Maybe post again.
Cheers,
Stan H.